Circular No. 8404

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004dk, 2004dw, 2004dy, AND 2004eg A. V. Filippenko, M. Ganeshalingam, F. J. D. Serduke, and J. L.

Hoffman, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Sept. 10 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2004dw (IAUC 8392) is of type II, perhaps 2 months past maximum brightness. The Hα P-Cyg profile is greatly dominated by the emission component, while the other hydrogen Balmer lines have typical P-Cyg profiles. SN 2004eg (IAUC 8401) is also of type II, but the velocity of the ejecta is extremely low, roughly 500 km/s, as measured from the absorption minima of the P-Cyg profiles. SN 2004dy (IAUC 8395) is a supernova, probably of type II, but the spectrum is rather noisy; relatively broad Hα emission seems to be present. SN 2004dk (IAUC 8377) was previously classified by Patat et al. (IAUC 8379) as a type-Ic supernova, but in the Lick spectrum, prominent He I lines have appeared, typical of type-Ib supernovae about 1 month past maximum brightness; indeed, the Lick series of three spectra obtained over the past few weeks shows an evolution more consistent with that of Ib than Ic supernovae.

NOVAE IN M31 K. Hornoch, Lelekovice, Czech Republic, reports the discovery of two

apparent novae in M31 by P. Kušnirák (first object), M. Wolf (second object), and himself (both objects) on CCD R-band images taken by them and L. Sarounova with the 0.65-m reflector at Ondřejov. The first new object is located at α = 0^{h}43^{m}26^{s}.84, δ = +41^{o}16′40′′.8 (equinox 2000.0), which is 479′′ east and 32′′ north of the center of M31. R-band magnitudes for this object from Hornoch: Aug. 2.974 UT, [20.5; 3.960, 19.2:; 4.901, 18.8; 6.997, 17.8; 7.982, 17.3; 9.070, 17.8; 10.005, 18.3; 10.956, 17.7; 12.057, 18.3; 15.943, 17.6; 23.902, 17.5; 24.060, 17.6; 28.899, 18.1; 31.915, 18.6; Sept. 3.018, 18.6; 3.868, 18.7; 4.819, 19.0; 5.058, 18.9; 6.930, 19.0; 8.929, 18.3; 9.893, 18.3; on Aug. 24.063, V = 17.6. The second new object is located at α = 0^{h}42^{m}40^{s}.25, δ = +41^{o}14′42′′.9, which is 46′′ west and 86′′ south of the center of M31; R-band magnitudes for this object from Hornoch: Aug. 24.060, [19.8; 28.899, [19.1; 31.915, 18.3; 31.938, 18.0; 31.952, 17.8; Sept. 3.018, 17.5; 3.868, 17.6; 4.841, 18.2; 5.862, 18.1; 6.930, 18.0; 8.953, 18.1; 9.893, 18.4. Filippenko et al. add that inspection of spectra, obtained at Lick (as described above) on Sept. 10, shows both objects to be novae with very strong Hα emission (FWHM ∼ 2000 km/s); considerably weaker Hβ and Fe II may be visible, as well.

2004 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green